You are here: GSI Wiki>FRS Web>S2VTPC21 (2026-10-30, NisanYaprakBenal)

S2 — VTPC21 (GTS3EG5T)

FRS Home Page - Areas.gif – S2 – Pictures – Drawings – Cable Connections – TPC Electronics – Memo for preparation before beam time

Updated November 07, 2019

Official Manual of TPC ( pdf.gif PDF )
Recent Log and Memo about TPC (2019) (External link)

Present Configuration (as of 2020/Oct/22)

- in a box (pocket) for the S2 vacuum chamber (GTS3EG7T)
- Window = 100 um Stainless steel (each of front and behind)
- Gap = 8 cm (drift HV = -3200 V)
- Normal orientation
- Now calibration positions should be x = -12, 0, +12 mm and y = -6, 0, +6 mm
- To have consistent track from all 4 TPCs, one could add the values on page 5 of this memo to the grid position
- Maximum allowed HV: -3200 V (drift), +1100 V (anode), -700 V (calibration, PMT).
For heavy ions, anode and PMT voltages should be much lower than the maximum values.

Status during previous beam time

2021/Feb/15-Feb/22 (12C beam, S533/BARB + EXPERT test)

- HV(anode) = +1050V, HV(drift) = -3200V, HV(grid-pmt) = -700V
- Analog raw signal around 150--200 mV. Vth(tpc module) around 45--60 mV
- Calibration file is /d/land4/sec/s533/tpc22_calib_0024.lmd (Logbook p41-43)
- Parameter sets used during experiment is found here.
2020/June/05--June/08 (proton beam, WASA000)

- Applied voltages: anode HV = +1150 V, drift = -3200 V, calib-pmt =
- Calibration file: wasa2020june_tpc22calib_[0015-0018].lmd
2020/May/xx--May/xx (Pb beam, S469, very low evenry)
2020/Apr/08--Apr/27 (Pb beam, S468, S482, S481)

- Applied voltages: anode HV = +580 V, drift HV = -3200V, fiber-PMT = -350V
2020/Feb/28--Mar/16 (Xe beam, S474, S459, S480)

- Applied voltages: anode HV = +600 V, drift HV = -3200V, fiber-PMT = -350V
- Used very short time to get the first impression about the beam poisition. Calibration grid (run0114). Parameters
- There are files with all TPC21+22+23+24, which can be used for overall check.
2020/Feb/20--Feb/26 (Kr beam, Ca fragments, Cave-C S467)

- Applied voltages: anode HV = +700 V, drift HV = -3200V.
- Used very short time to get the first impression about the beam poisition. No grid caliberation was performed.
- There are files with all TPC21+22+23+24, which can be used for calibration (if needed). Runs 92--95.
2019/Dec/09--Dec/19 (U beam)

- After changing primary beam from 40Ar to 238U.
- Applied voltages: anode HV = +550 V, drift = -3200V
2019/Nov/18--Dec/06 (Ar beam)

- Installed at S2 as TPC22.
- All signals were checked and ok, calibration scintllator double peak ?? >> to be investigated.
- Gap is 8cm. normal orientation. DL1 signal is fine (see comment in 2019-April run).
- Applied voltages: anode HV = +800 V, drift = -3200V, fiber PMT = -500V.
2019/April/1--3 (107Ag, 290MeV/u, primary beam)

- Installed at S2 as TPC21.
- Signal of TPC22-DL1 is abnormally small. Swapping cables did not work. Something in the TPC chamber should be the reason.
- Other signals are reasonable.
- Calibration scintllator was ok.
- normal orientation - We need to open this TPC after the beam time to investigate problems of DL1.

On 28/11/2023 the thresholds of the TPC raw signal were measured and the measurements are reported in Table 1.

Table 1. TPC thresholds on the discriminator/amplifier

Signals TPC21 TPC22 TPC23 TPC24 TPC31 TPC41 TPC42
A11 no access -73,5 no access -144,4 -366 -154,3 -163,8
A12 -72,7 no access -159,3 -147,6 -387 -144,9 -164,9
A21 no access -77 -158,3 -159,9 -353 -147,8 -160,7
A22 -157,1 -76,4 -144,7 -156,1 -405 -150,1 -166,8
DL1 -112,2 -108,9 -127,3 -195,4 -83,3 -152,2 -119,9
DR1 -79,7 -77,1 -117,3 -207 -110,9 -154,7 -116,4
DL2 -184,4 -71,3 -122,7 -196,8 -64,3 -150,6 -120,9
DR2 -180 -71 no access -190,8 -49,6 -148,2 -127
On 28/11/2023 the thresholds of the TPC Sci-Sig raw were measured and the measurements are reported in Table 2.

Table 2. TPC calibration grid values.

TPC Threshold on discriminator High Voltage (V) Current at no bema (\mu A)
TPC21 -64,5 600,5 214
TPC22 -46,5 600 213,5
TPC23 -136,5 500,5 178
TPC24 -15,3 500,5 178
TPC31   330 117,5
TPC41 -49,6    
TPC42 -45,6    
-- NisanYaprakBenal - 2026-10-27

On 28/11/2023 the readout window of the TDC for the TPC has been moved about 1 microsecond in order to better center the events.

-- SagarRoy - 2021-06-26
I Attachment Action Size Date Who Comment
Areas.gifgif Areas.gif manage 1 K 2021-07-10 - 11:25 SagarRoy  
pdf.gifgif pdf.gif manage 128 bytes 2021-07-14 - 20:06 SagarRoy  
tpcmanual.pdfpdf tpcmanual.pdf manage 2 MB 2021-07-14 - 19:58 SagarRoy  
Please login to edit this topic
Topic revision: r8 - 2026-10-30, NisanYaprakBenal
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding GSI Wiki? Send feedback | Legal notice | Privacy Policy (german)